OpenStudy (anonymous):

]Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1/4x + 2. y = 4x - 1 y = 4x - 3 y =1/4 x - 3 y =1/4 x - 1

OpenStudy (anonymous):

We work with the slope-intercept form of the line: y = mx + b where "m" is the slope and "b" is the y-intercept. Since we want a parallel line, we will have the same slope of m = 1/4. Now, we use the point-slope form:\[y - y _{1} = m(x - x _{1})\] y - (-2) = m(x - (-4)) y + 2 = (1/4)x + 1 y = (1/4)x - 1 the last selection.

OpenStudy (anonymous):

i need help with only two more questions, would you like to help?

OpenStudy (anonymous):

Sure! Fire away!

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2 y = 1/2x + 5 y =1/2 x - 3 y =1/2 x - 5 y =1/2 x + 3

OpenStudy (anonymous):

and ]Choose the equation of the line passing through the point (-1, 7) and perpendicular to y = -1/4x - 5. y = -4x - 11 y = -4x - 3 y = 4x + 11 y = 4x + 3

OpenStudy (anonymous):

For the first of these 2. Similar to the last problem, using the point-slope form: y - (-4) = (1/2)(x - 2) y + 4 = (1/2)x - 1 y = (1/2)x - 5

OpenStudy (anonymous):

For the second of these last 2, Similar, but here we have to find the slope. We are being asked to get a perpendicular line which will have a negative reciprocal for the slope or "4" because (4)(-1/4) = -1. y - 7 = 4(x - (-1)) y - 7 = 4x + 4 y = 4x + 11

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -5 and a y-intercept of -1.

OpenStudy (anonymous):

You're the best :D ♥♥♥

OpenStudy (anonymous):

You're entirely welcome!